# Spatiotemporal Semantic V2X Framework for Cooperative Collision Prediction

### v1

@@ IV-B Numerical Results
In this part, performance evaluation and numerical results for both the communication and the V-JEPA model are analyzed. In the communication part, there are two criteria: average compression ratio and transmission latency, as shown in Figure4. Note that we set SnR for BPSK and QAM16 to 12 dB and 22 dB, which are proven to be most efficient values in our previous work [12]
Figure 4: Results for Average Compression Ratio and Average V2X Latency For Each Embedding Format
The above part of Figure4 shows the compression ratio, while the bottom plot illustrates the average V2X transmission ratio. Raw videos collected in various experiments have variable durations, impacting their total size. In terms of compression ratio, we categorized videos into lower, standard, and higher duration categories. We calculated the compression ratios for each category’s embedding encodings (FP32, FP16, INT8), as explained in III-D.
In transmission latency, for BPSK and QAM16 modulation schemes, we set uplink data rate to 10 MHz and 20 MHz, and calculate the average transmission latency for FP32, FP16, and INT8 embedding encodings. Once the uplink rate increases, average transmission latency is reduced for both modulation schemes, and the results prove that QAM16 has much lower latency for BPKS, which makes it more suitable for real-time.
On the other hand, the evaluation of the V-JEPA model on the traffic collision avoidance is performed in two steps. In the first step, we determine which post-processing methods (Section III-A2) yield the optimal performance for V-JEPA model collision estimation. Then, we select the best postprocessing and adjust the interval before the collision occurs to see how early and accurately it can be detected. Results for different post-processing methods and original clips can be seen in Figure5 via confusion matrices.
@@
Figure5 presents 4 confusion matrices for V-JEPA models; No post-processing (upper left), heatmap (upper right), binary masked (bottom left), and heatmap+binary masked (bottom right). For collision scenarios, each video clip was trimmed to exclude 8 frames depicting or directly preceding the collision event. Model achieves 76% F1 scores without post-processing, 80% heatmap, 84% binary masked, and 81.81% heatmap+binary masked. Heatmap+binary masked (3) has a lower false positive than single binary masked (6), but binary masked has the highest F1 scores due to its low false negatives. This indicates that YOLO-based vehicle highlighting compromises performance in false negative cases, despite most collisions being detected properly. In case 2, we examine how the number of removed frames before collisions affects performance (see Table II).
TABLE II: Results for Different Previous-Frames Gap Before Collisions
@@ V Conclusions
This paper presented a V2X SemComm framework for proactive collision prediction using the V-JEPA. RSUs serve as semantic encoders, generating predictive spatiotemporal embeddings from video data captured via the QLabs digital twin. Post-processed clips highlighting critical regions were used with pretrained V-JEPA encoders and K400 attentive probes to convert frame sequences into compact vectors capturing essential motion cues, achieving an 84% F1 score. By transmitting low-dimensional embeddings instead of raw video, the framework reduces communication payloads by up to four orders of magnitude, enabling bandwidth-efficient and low-latency semantic message exchange. Future work will explore adaptive semantic compression and cross-vehicle embedding fusion to enhance cooperative perception and scalability in next-generation V2X systems.
@@ Acknowledgment
This work is supported in part by the MITACS Accelerate Program, in part by the NSERC CREATE TRAVERSAL program, and in part by the Ontario Research Fund-Research Excellence program under grant number ORF-RE012-026. The authors would like to thank Quanser for their support in the generation of the traffic data via Quanser Interactive Lab (https://www.quanser.com/products/quanser-interactive-labs/).

### v2

@@ IV-B Numerical Results
From a communication perspective, two key metrics are analyzed: compression ratio and Tx latency (Figure 4).
Figure 4: Average compression ratio for different video lengths and Tx latency for different modulation schemes
The left portion of Figure 4 illustrates the achieved compression ratios, while the right portion depicts the corresponding Tx latencies. The raw videos obtained from different experiments vary in duration, resulting in different overall data sizes. To assess compression efficiency, the videos are categorized into two duration groups: normal-length and large-length clips. For each group, compression ratios are computed across the three encoding formats (FP32, FP16, and INT8), as described in Section III-D. In both cases, INT8 achieves the highest compression ratio, exceeding $10^{5}$, followed by FP16 and FP32. Large-length clips (64 frames) yield higher ratios than normal-length clips (33 frames), reflecting reduced transmission payload with longer videos. For example, assuming $N=64$, $H_{o}=W_{o}=2048$, the raw video size is approximately $64\times 2048\times 2048\times 3=0.81$ GB, while the transmitted semantic embeddings require only $1\times 1280\times b=0.005$ MB, $0.0025$ MB, and $0.0012$ MB for FP32, FP16, and INT8, respectively, corresponding to compression gains of $1.6\times 10^{5}$, $3.2\times 10^{5}$, and $6.4\times 10^{5}$.
To evaluate latency and Tx data rate, the SNR values for BPSK and QAM16 are set to 12 dB and 22 dB, respectively, as identified optimal previously [17]. QAM16 achieves the lowest Tx latency across all encodings, 0.27, 0.13, and 0.06 ms for FP32, FP16, and INT8,while BPSK yields 0.50, 0.25, and 0.12 ms under the same conditions. All values remain well below the 5 ms V2X threshold, confirming the framework’s suitability for real-time, low-latency vehicular communication.
The evaluation of the V-JEPA model for traffic collision avoidance is conducted in two stages. First, various post-processing methods (Section III-A2) are analyzed to identify the configuration that provides the highest collision prediction accuracy. Next, the selected post-processing approach is used to adjust the frame gap preceding the collision to assess how early and reliably potential collisions can be detected. The resulting confusion matrices for different post-processing methods and the original clips are presented in Figure 5.
@@
Figure5 presents the confusion matrices for no post-processing (upper left), heatmap (upper right), binary masked (bottom left), and heatmap+binary masked (bottom right). For collision scenarios, each video clip was trimmed to exclude 8 frames depicting or directly preceding the collision event. Model achieves 76% F1 scores without post-processing, 80% heatmap, 84% binary masked, and 81.81% heatmap+binary masked. Heatmap+binary masked has a lower false positive than binary masked, but binary masked has the highest F1 scores due to its low false negatives. This indicates that YOLO-based vehicle highlighting compromises performance in false negative cases, despite most collisions being detected properly. In the second step, we examine how the number of removed frames (frame-gap) before collisions affects collision prediction performance of V-JEPA Model (see Table II).
TABLE II: Results for Different Previous-Frames Gap
@@
The number of removed frames before the collisions indicates how early we can detect the collision. If the frame-gap before collision increases, the model tries to estimate the likelihood of future collision earlier. To achieve better results, we set the frame-gap to 12 (left), 8 (middle), and 4 (right) with binary mapped clips, resulting in 73.68%, 84%, and 79% F1 scores, respectively. With 12 frames removed before the collision, the model yields highly false negative values because many collisions occur during cornering, and a longer frame cut results in clips ending before cornering begins. The 4-frame gap has a slightly lower F1 score than 8-frame gap before colliding because the model is already cornering and 4-frame removing used longer clips, making the model’s work challenging.
@@ V Conclusions
This paper has presented a V2X SemComm framework for proactive collision prediction using the V-JEPA. RSUs serve as semantic encoders, generating predictive spatiotemporal embeddings from video data captured via the QLabs digital twin. Post-processed clips highlighting critical regions were used with pretrained V-JEPA encoders and K400 attentive probes to convert frame sequences into compact vectors capturing essential motion cues, achieving 92% accuracy and 8% F1-score improvement in collision prediction. By transmitting low-dimensional embeddings instead of raw video, the framework reduces communication payloads by up to five orders of magnitude, enabling bandwidth-efficient and low-latency semantic message exchange. We are currently exploring cross-vehicle embedding fusion to enhance cooperative perception and scalability in next-generation V2X systems while working on the real dataset in more complex scenarios.
@@ Acknowledgment
This work is supported in part by the MITACS Accelerate Program, in part by the NSERC CREATE TRAVERSAL program, and in part by the Ontario Research Fund-Research Excellence program under grant number ORF-RE012-026. The authors would like to thank Quanser for their support in the generation of the traffic data via Quanser Interactive Lab.